mikewwe14:

Which formula shows how the length of a rectangle can be determined from the permieter and the width?

mikewwe14:

@Shadow

mikewwe14:

@Vocaloid

mikewwe13:

1 attachment
Shadow:

There we go ._.

Vocaloid:

P = 2L + 2W solve for L

mikewwe13:

l = -w + P/2

mikewwe13:

am i correct ?

Vocaloid:

the math is correct but you may want to re-write in a form that looks more like the answer choices for example: P = 2L + 2W subtracting 2W from both sides gives us (P - 2W) = 2L then dividing by 2 gives us (P-2W)/2 = L which answer choice corresponds with this formula?

Vocaloid:

notice how W is in the numerator not the denominator

mikewwe13:

B

Vocaloid:

good